# Unsupervised machine learning for the detection of exotic phases in skyrmion phase diagrams

### v1

@@ Abstract
Undoubtedly, machine learning (ML) techniques are being increasingly applied to a wide range of situations in the field of condensed matter. Amongst these techniques, unsupervised techniques are especially attractive, since they imply the possibility of extracting information from the data without previous labeling. In this work, we resort to the technique known as “anomaly detection” to explore potential exotic phases in skyrmion phase diagrams, using two different algorithms: Principal Component Analysis (PCA) and a Convolutional Autoencoder (CAE). First, we train these algorithms with an artificial dataset of skyrmion lattices constructed from an analytical parametrization, for different magnetizations, skyrmion lattice orientations, and skyrmion radii. We apply the trained algorithms to a set of snapshots obtained from Monte Carlo simulations for three ferromagnetic skyrmion models: two including in-plane Dzyaloshinskii-Moriya interaction (DMI) in the triangular and kagome lattices, and one with an additional out of plane DMI in the kagome lattice. Then, we compare the root mean square error (RMSE) and the binary cross entropy (BCE) between the input and output snapshots as a function of the external magnetic field and temperature. We find that the RMSE error and its variance in the CAE case may be useful to not only detect exotic low temperature phases, but also to differentiate between the characteristic low temperature orderings of a skyrmion phase diagram (helical, skyrmions and ferromagnetic order). Moreover, we apply the skyrmion trained CAE to two antiferromagnetic models in the triangular lattice, one that gives rise to antiferromagnetic skyrmions, and the pure exchange antiferromagnetic case. Despite the predictably larger RMSE, we find that, even in these cases, the RMSE is also an indicator of different orderings and the emergence of particular features, such as the well-known pseudo-plateau in the pure exchange case.
@@ I Introduction
Machine Learning (ML) has been increasingly incorporated into a wide variety of technological and scientific research in the last few years. In condensed matter physics, some of the first studies using ML include the representation of quantum states [1], the identification of phases of matter [2], inverse design in photonics [3] and the study of phase transitions [4, 5, 6]. In particular, ML has been used to explore different aspects of magnetic skyrmions, which are topological textures [7, 8] with potential technological applications, especially in memory storage devices [9, 10]. Feed forward and convolutional neural network have been used to classify different magnetic phases and predict features in simple skyrmion models using simulations data [11, 12, 13, 14, 15, 16, 17, 18]. Moreover, ML algorithms have been applied also to experimental data to explore skyrmion phases [19, 20, 21].
In general, ML techniques could be devided in three large types of techniques [22, 23]: supervised, unsupervised and reinforcement learning, the latter one being the case where the algorithm “learns” solely from the data. In supervised ML, a previous labeling and knowledge of the data is needed before applying the algorithms to an unknown dataset, and is widely used in classification and regression tasks. On the other hand, no data labeling is required in unsupervised ML, and some of these techniques may be used for example to group similar instances of the data (clustering) or detect atypical instances (anomaly detection). A variety of this type of unsupervised (or semi-supervised) methods has been used in condensed matter physics, from Principal Component Analysis (PCA) in Ising systems [5] to support vector machines with tensorial kernels applied to models in different types of topology [24, 25, 26, 27, 28, 29, 30, 31], the use of autoencoders to study neutron scattering data in spin ice systems [32, 33] and to explore anomaly detection in simple frustrated models [34].
@@ II Brief overview on skyrmions and spin models
Magnetic skyrmions are an arrangement of spins characterised by the topological charge, defined as ; are the spin unit vectors, with promising technological applications due to their robustness, stability and easy manipulation. A large number of materials and models have been shown to host skyrmions and skyrmion like textures (such as antiferromagnetic skyrmions [36, 37], magnetic bubbles, antiskyrmions [38, 39, 40], merons [41], etc). A skyrmion crystal or lattice in a bidimensional system is a periodic arrangement of skyrmions and may simply parametrised as the superposition of three non-coplanar helices [42, 7]; the spin parametrization at position is given by :
@@
The structure factor, which is the Fourier transform of the spin-spin correlations, of this type of arrangement gives six bright peaks, corresponding to the three inequivalent ordering vectors. Thus, a “triple-q” pattern in reciprocal space is a strong suggestion that a skyrmion lattice is present, as measured first in bulk MnSi [43]. Another parameter that may be calculated in a discrete lattice is the scalar chirality, the sum of the triple product of neighbouring spins () forming a triangle . This quantity is the discrete version of the topological charge, which is for skyrmions [7]. Therefore, for a given arrangement of spins, a non zero value of the chirality may be an indicator of skyrmions, regardless of whether they are organized periodically or not.
@@
As was introduced in [44], a possible simple model to realise these periodic skyrmion arrangements in a bidimensional lattice is one where ferromagnetic nearest neighbors coupling competes with antisymmetric in-plane Dzyaloshinskii Moriya (DM) interactions under an external magnetic field:
here the magnetic moment or spin is a three dimensional unit vector (Heisenberg spins) at site , is the exchange coupling, the in-plane DMI along the bonds of the chosen lattice (, with absolute value ), and an external magnetic field perpendicular to the lattice plane. At low temperatures and no external field a helical order is found. Under an increasing external magnetic field, a skyrmion crystal is stabilized, and then at higher fields all spins are aligned with the field giving rise to a ferromagnetic or field polarized phase. Intermediate phases are enhanced with temperature: a bimeron phase at lower fields, as helices are distorted into skyrmions, and a skyrmion gas at higher magnetizations, where skyrmions are stabilized in a non-periodic arrangement [44, 45]. The inclusion of additional interactions may imply deviations from this simple scheme. For example, it has been shown that an increasing additional in plane site anistropy may take the skyrmion lattice to an arrangement of merons, which are topological structures with half the skyrmion topological charge [41], and more complex phase diagrams may also be found in models with frustrating exchange interactions and no DMI [46, 47].
@@
Since we aim to explore techniques that may detect exotic phases, we focus on three models. First, we take the Hamiltonian from Eq. (2) in two triangular based lattices: triangular and kagome, which may itself be devided into three triangular lattices formed by the three types of spins in each elementary plaquette. Then, as a third model we take the Hamiltonian from Eq. (2) in the kagome lattice and include an out-of-plane DMI () at the specific value . The combination of this out-of-plane and in plane gives rise to several exotic phenomena, most notably a stabilization of a skyrmion gas at higher temperatures, which preceds the formation of a skyrmion crystal, and a “bimeron glass” phase at higher magnetic fields and lower temperatures, between the skyrmion crystal and the fully polarized phase [35, 49]. Here, we will show that it may be possible to detect these phases using unsupervised machine learning techniques, without inspection of the snapshots.
@@
In Table 1 we list these three different spin models. We label the one in the triangular with ferromagnetic exchange coupling and in-plane DMI along the bonds TFDMxy. In analogy, the same model but in the kagome lattice is named KFDMxy. Finally, the KFDMz spin model is as the KFDMxy with additional out of plane DM, . Notice that is fixed to in all three spin models, but the is larger for the triangular lattice. This is simply so that the radius of the resulting skyrmions be similar to that of the KF spin models.
@@ III Machine Learning scheme and dataset
We study the behaviour of the RMSE and BCE for different parameters and comparing models, to evaluate whether we may obtain information on the physics of each model, and explore the possibility of using this technique to find exotic phases in skyrmion-hosting systems.
@@ III.1 Training dataset
We construct “artificial” skyrmions cristals in triangular lattices with sites, following the parametrization presented in Eq. (1). We take 7 possible values for the skyrmion radii and 10 values of the parameter , which after normalization changes the magnetization of the lattice. Since in our analysis we will be using snapshots from MC simulations as images, in this constructed dataset we do not require periodic boundary conditions, but we must take into account rotations and translations of the parametrization pattern, which we do considering 8 possible rotations and 3 different translations in the position along the 3 directions of the nearest neighbors bonds of the triangular lattice. Therefore, we have a training and validation dataset of different skyrmion crystals in triangular lattices.
@@ III.2 Monte Carlo simulations dataset
We performed Monte Carlo Metropolis-Hastings simulations, combined with microcanonical updated (overrelaxation) on the three spin models described in the previous section: (1) TFDMxy: a ferromagnetic exchange and in plane DMI model in a triangular lattice under a magnetic field, fixing (2) KFDMxy: a similar model in a kagome lattice, for (3) KFDMz: the model presented in Eq. (2) including an out of plane Dzyaloshinskii-Moriya interaction . Simulations were done using periodic boundary conditions, and lowering the temperature from high temperature in 80 steps using the annealing technique. Up to 10 copies with independent seeds were done for each model and magnetic field. To apply the algorithms to the kagome snapshots, we chose one of the three sublattices (each one corresponding to the three types of sites in the kagome lattice, numbered in Table 1), so as to work with a snapshot.
@@ III.3 Algorithms
Here we give details on the two algorithms used in this work: Principal Component Analysis (PCA) and Convolutional Auto Encoder (CAE).
@@ III.3.2 Convolutional Autoencoder
In our work, we will use the “anomaly detection” approach. The idea is to train an AE for a given type of data, and then apply it to a larger dataset, comparing the input and the output with a convenient variable, such as the root mean square error. If the error is small, then the input and output are similar, if it is large, then the input data may not be of the same type as that used in training the AE, so a larger error signals an anomaly. As in the PCA, we will use the artificial skyrmion lattice dataset to train a convolutional autoencoder (CAE - an AE with a convolutional layer), measuring the RMSE. In this case, 80 of the data were used for training and 20 for validation. The RMSE for the trained CAE was 0.00071 in the training set and 0.00073 in the validation set.
The structure of the CAE is the following: the encoder consists of one convolutional layer with activation function ‘relu’, filter size 3 and padding ‘same’, followed by a MaxPooling layer with filter size 2 and padding ‘same’. The decoder consists of a convolutional layer, with the same carachteristics as the first one, and an UpSampling layer. It finalizes with a convolutional layer with one filter of size 3, a sigmoid activation function and padding ‘same’. The CAE was trained and validated with batch size 64, using early stopping for a maximum of 300 epochs with a patience of 10. Implementation of the CAE was done in TensorFlow using Keras [55, 56].
@@ IV Results and discussion
In this section, we present the analysis of the comparison between the input and output data for the MC snapshots obtained when applying the PCA and CAE algorithms that were trained with the artificial set of triple-q skyrmion lattices. We start in Subsec. A focusing on the three ferromagnetic models introduced in Sec. II (Table 1). Then, we will explore whether the anomaly detection technique using the same skyrmion-lattice trained CAE is able to provide some insight in two different antiferromagnetic models, which we will describe in Subsec. B.
@@ IV.1 Application to ferromagnetic skyrmion models
First, we compare the RMSE obtained for both algorithms for the TFDMxy snapshots. In order for the method to work, we would expect the RMSE to be lower at the skyrmion lattice phase, which is found at intermediate magnetic fields and lower temperature, and higher in the other phases, such as the low field helical phase at lower magnetic fields, or the high temperature paramagnetic phase. In Fig. 1, we present a series of examples comparing the input Monte Carlo snapshot with the output (decoded) snapshot after applying trained PCA and CAE. The first four rows correspond to the low temperature snapshots, for four phases that arise as the magnetic field is increased: helical, skyrmion lattice, skyrmion gas and ferromagnetic. At first glance, the low temperature decoded images seem very close to the input ones, both for PCA and CAE, even for the helical and the ferromagnetic cases, which are types of configurations that were not used to train the algorithms. In these examples, the RMSE (indicated in each decoded image) for PCA is even lower in the helical and ferromagnetic cases than in the skyrmion lattice. This is not ideal, since we are proposing to use a larger RMSE as an indicator that the system is further from an ordered skyrmion phase. However, there is a significant difference in the output images at higher temperatures, illustrated in the last row of Fig 1 for a snapshot in the paramagnetic phase: both the PCA and the CAE have a significantly higher RMSE.
@@
To further explore and quantify whether the RMSE obtained with PCA or CAE is a useful tool to distinguish between the low temperature phases, in Fig. 2 we present these RMSE as a function of the external magnetic field at the lowest simulated temperature for the TFDMxy model (). The curves are averaged over 10 independent copies, and the error bars are calculated as the standard deviation. The background color in Fig. 2 are a guide to the eye and indicate three type of phases: a helical phase, a skyrmion phase (which includes skyrmion lattices and skyrmion gas) and a ferromagnetic or field polarized region. We also present the absolute value of the scalar chirality (scaled by a factor for better comparison), the MC calculated parameter that distinguishes between the different phases: it is zero at low field in the helical case, goes up significantly where the skyrmion lattices are stabilized, gets lower as these lattices go into a skyrmion gas with increasing magnetic field, and finally is zero when all spins are aligned with the external field. Let us recall that the chirality cannot be “learned” by the algorithms, since in this proposal only the spin component is used for training.
@@
As for the RMSE calculated from the output images of both algorithms, it can be seen that the RMSE from PCA is in fact lowest for the ferromagnetic phase, highest for the skyrmion gas, and, within the error, does not significantly distinguish between the helical and skyrmion lattice phases. On the other hand, the RMSE obtained comparing the MC snapshots and the CAE decoded ones shows a more interesting behaviour, clearly distinguishing three regions: a low field one with higher RMSE, corresponding to the helical phase, an intermediate one with the lowest values (the skyrmion phases), and then a third region where the RMSE goes slightly but noticeably up with and flattens at the ferromagnetic phase. We also present the binary cross entropy curve for the CAE. The BCE (scaled by a factor for better comparison) also seems to distinguish between these three regions, showing an abrupt change as the system enters the skyrmion phases, but it is lowest in the ferromagnetic phase, and not in the skyrmion region, where the CAE was trained. Therefore, we choose the RMSE from the CAE to apply the anomaly detection technique in other models. As a comment, a closer inspection of the RMSE curve also suggests that the RMSE error may also be a useful variable, an idea we will shortly come back to.
@@
Having chosen our variable, the RMSE between the MC input snapshot and the decoded output from the trained CAE, we now apply the CAE to the other two ferromagnetic skyrmion models, both in the kagome lattice with in-plane interactions (KFDMxy), and one with an additional out of plane DMI (KFDMz) (see Table 1). We plot the RMSE and show the MC chirality 111We remind the reader that here we are choosing one kagome sublattice to apply the CAE, since it is trained for triangulat lattices. Thus, the chirality presented here is the sublattice chirality. and magnetization as a function of magnetic field at the lowest simulated temperature for both models in Fig. 3, where all curves are normalised to 1 to qualitatevely compare the behaviours. For the KFDMxy, left panel, the behavior is very similar to the triangular lattice case. However, for the KFDMz model, before flattening in the ferromagnetic phase, the RMSE goes clearly up. In the MC variables this is connected to a jump in the magnetization, but no significant differences are seen in the chirality, which is the order parameter connected to the skyrmion phases. Thus, the RMSE suggests that there is a different phase at higher magnetic fields, for . Inspection of the snapshots reveals this is indeed the case; an example is shown in Fig. 4, top row. It is an unusual high field bimeron + skyrmion phase, with possible glassy behaviour [35]. Regardless of the nature of the ptextures, the RMSE indicates that at higher fields the KFDMz presents a different phase than the other two ferromagnetic models. We compare the mean value and the standard deviation of the RMSE as a function of the external field for the three models in Fig. 5. In the plot, is scaled for the TFDMz model, but the RMSE is not. From the compariston, it is clear that the RMSE is quantitavely and qualitatevely very similar in the three models in three regions (corresponding to helices, skyrmions and ferromagnetic), but evidently there is an exotic phase for the KFDMz case before the system goes into the ferromagnetic phase. The jump of the RMSE for this model in this region is not seen in the other two cases, strongly supporting the idea that in that region of the system goes into a different type of ordering.

### v2

@@ Abstract
Undoubtedly, machine learning (ML) techniques are being increasingly applied to a wide range of situations in the field of condensed matter. Amongst these techniques, unsupervised techniques are especially atractive, since they imply the possibility of extracting information from the data without previous labeling. In this work, we resort to the technique known as “anomaly detection” to explore potential exotic phases in skyrmion phase diagrams, using two different algorithms: Principal Component Analysis (PCA) and a Convolutional Autoencoder (CAE). First, we train these algorithms with an artificial dataset of skyrmion lattices constructed from an analytical parametrization, for different magnetizations, skyrmion lattice orientations, and skyrmion radii. We apply the trained algorithms to a set of snapshots obtained from Monte Carlo simulations for three ferromagnetic skyrmion models: two including in-plane Dzyaloshinskii-Moriya interaction (DMI) in the triangular and kagome lattices, and one with an additional out-of-plane DMI in the kagome lattice. Then, we compare the root mean square error (RMSE) and the binary cross entropy (BCE) between the input and output snapshots as a function of the external magnetic field and temperature. We find that the RMSE error and its variance in the CAE case may be useful to not only detect exotic low temperature phases, but also to differentiate between the characteristic low temperature orderings of a skyrmion phase diagram (helical, skyrmions and ferromagetic order). Moreover, we apply the skyrmion trained CAE to two antiferromagnetic models in the triangular lattice, one that gives rise to antiferromagnetic skyrmions, and the pure exchange antiferromagnetic case, finding that the behaviour of the RMSE is still an indicator of different ordering. Finally, we explore the portability of the technique with the same CAE applying it to a ferromagnetic skyrmion model in the square lattice.
@@ I Introduction
Machine Learning (ML) has been increasingly incorporated into a wide variety of technological and scientific research in the last few years. In condensed matter physics, some of the first studies using ML include the representation of quantum states [1], the identification of phases of matter [2], inverse design in photonics [3] and the study of phase transitions [4, 5, 6]. In particular, ML has been used to explore different aspects of magnetic skyrmions, which are topological textures [7, 8, 9, 10, 11] with potential technological applications, especially in memory storage devices [12, 13, 14]. Feed forward and convolutional neural network have been used to classify different magnetic phases and predict features in simple skyrmion models using simulations data [15, 16, 17, 18, 19, 20, 21, 22]. Moreover, ML algorithms have been applied also to experimental data to explore skyrmion phases [23, 24, 25].
In general, ML techniques could be devided in three large types of techniques [26, 27]: supervised, unsupervised and reinforcement learning, the latter one being the case where the algorithm “learns” solely from the data. In supervised ML, a previous labeling and knowledge of the data is needed before applying the algorithms to an unknown dataset, and is widely used in classification and regression tasks. On the other hand, no data labeling is required in unsupervised ML, and some of these techniques may be used for example to group similar instances of the data (clustering) or detect atypical instances (anomaly detection). A variety of this type of unsupervised (or semi-supervised) methods has been used in condensed matter physics, from Principal Component Analysis (PCA) in Ising systems [5] to support vector machines with tensorial kernels applied to models in different types of topology [28, 29, 30, 31, 32, 33, 34, 35], the use of autoencoders to study neutron scattering data in spin ice systems [36, 37], the identification of superconducting phases [38] and to explore anomaly detection in simple frustrated models [39].
@@ II Brief overview on skyrmions and spin models
Magnetic skyrmions are an arrangement of spins characterised by the topological charge, defined as ; where are the spin unit vectors. Skyrmions are especially suitable for promising technological applications due to their robustness, stability and easy manipulation. A large number of materials and models have been shown to host skyrmions [41, 42] and skyrmion like textures (such as antiferromagnetic skyrmions [43, 44], magnetic bubbles, antiskyrmions [45, 46, 47], merons [48], etc). A skyrmion crystal or lattice in a bidimensional system is a periodic arrangement of skyrmions and may simply be parametrised as the superposition of three non-coplanar helices [49, 10]; the spin parametrization at position is given by :
@@
A relevant parameter that may be calculated in a discrete lattice of spins is the scalar chirality, the sum of the triple product of neighbouring spins () forming a triangle, defined as:
This quantity, with a factor of , is the discrete version of the topological charge, which is for skyrmions [10], and is related to the skyrmion number [50]. Therefore, for a given arrangement of spins, a non zero value of the chirality may be an indicator of skyrmions, regardless of whether they are organized periodically or not.
@@
As was introduced in [42], a possible simple model to realise these periodic skyrmion arrangements in a bidimensional lattice is one where ferromagnetic nearest neighbors coupling competes with antisymmetric in-plane Dzyaloshinskii Moriya interactions (DMI) [51, 52] under an external magnetic field:
here the magnetic moment or spin is a three dimensional unit vector (Heisenberg spins) at site , is the ferromagnetic exchange coupling, the in-plane DMI along the bonds of the chosen lattice (, with absolute value ), and an external magnetic field perpendicular to the lattice plane. In this model, the skyrmion radius is related to the ratio, a larger ratio implies smaller skyrmions.
We present a typical phase diagram for the model in Eq.(3) , in Fig. 2, top panel, constructed with the scalar chirality as a function of temperature and magnetic field , obtained from Monte Carlo simulations in a triangular lattice with and sites (). At low temperatures and no external field a helical order is found. Under an increasing external magnetic field, a skyrmion lattice is stabilized, and then at higher fields all spins are aligned with the field giving rise to a ferromagnetic or field polarized phase. Intermediate phases are enhanced with temperature: a bimeron phase at lower fields, between helices and the skyrmion lattice, as helices are broken and distorted into bimerons; and a skyrmion gas at higher magnetizations, between the skyrmion lattice and the ferromagnetic ordering, where the skyrmion lattice order is lost and a smaller number of skyrmions are found in a non-periodic arrangement as the system is further magnetized [42, 53]. In this work, the bimerons look as “elongated skyrmions”, and are actually a pair of merons (topological structures with ), see for example [53, 54], as opposed to meron - antimeron pairs as defined in [55]. Typical configurations of these phases are shown in Fig. 3.
In the bottom panel of Fig. 2 we focus on the low temperature phases, and compare the absolute value of the scalar chirality density with the magnetisation . We see that there are indeed three main phases: at low field there are helices with no chirality; at larger fields the chirality dramatically increases (and the magnetisation shows a sharp change), indicating the presence of skyrmions. As the field is increased, the number of skyrmions is reduced, reflected in a smaller , and then the system goes into the ferromagnetic phase (), where the chirality is null. The background colors are a guide to the eye and indicate three type of phases: a helical phase, a skyrmion phase (which includes skyrmion lattices and a small skyrmion gas region) and a ferromagnetic or field polarized order.
The inclusion of additional interactions may imply deviations from this simple scheme. For example, it has been shown that an increasing additional in plane site anistropy may take the skyrmion lattice to an arrangement of merons [48], and more complex phase diagrams may also be found in models with frustrating exchange interactions and no DMI [56, 57].
@@
Since we aim to explore techniques that may detect exotic phases, we focus on three models. First, we take the Hamiltonian from Eq. (3) in two triangular based lattices: triangular and kagome, which may itself be divided into three triangular lattices formed by the three types of spins in each elementary plaquette. Then, as a third model we take the Hamiltonian from Eq. (3) in the kagome lattice and include an out-of-plane DMI () at the specific value . It has been shown that in the ferromagnetic kagome lattice with out-of-plane DMI () a chiral spin liquid emerges, where the in-plane spin components retain the degeneracy and correlations of an algebraic classical spin liquid, and the system has a net magnetization [58]. The combination of this out-of-plane and in plane gives rise to several exotic phenomena, most notably a stabilization of a skyrmion gas with well defined skyrmion structures at higher temperatures, which precedes the formation of a skyrmion lattice, and a “bimeron glass” phase at higher magnetic fields and lower temperatures, between the skyrmion crystal and the fully polarized phase [40, 59], where the system retains chiral spin liquid behavior in the plane. This fully polarized phase extends at lower magnetic fields and intermediate temperatures, and allows the control of the number of skyrmions with temperature, going from a skyrmion gas to a skyrmion lattice as the temperature is lowered, which is not the typical way that skyrmion lattices are stabilized in two dimensional systems [60, 61, 62]. Here, we will show that it may be possible to detect these phases using unsupervised machine learning techniques, without inspection of the snapshots.
@@
In Table 1 we list these three different spin models. We label the one in the triangular lattice with ferromagnetic exchange coupling and in-plane DMI along the bonds TFDMxy. In analogy, the same model but in the kagome lattice is named KFDMxy. Finally, the KFDMz spin model is as the KFDMxy with additional out-of-plane DMI, . Unless otherwise indicated, we will refer as TFDMxy as the model in the triangular lattice with , where the radius of the resulting skyrmions be similar to that of the KF spin models. Nonetheless, we will also test our analysis comparing results for other values of the DMI in the triangular lattice.
@@ III Machine Learning scheme and dataset
We study the behaviour of the RMSE and BCE for different parameters and comparing models, to evaluate whether we may obtain information on the physics of each model, and explore the possibility of using the anomaly technique comparing the errors (RMSE and BCE) to find exotic phases in skyrmion-hosting systems.
@@ III.1 Training dataset
We construct “artificial” skyrmions cristals in triangular lattices with () sites, following the parametrization presented in Eq. (1). We take 7 possible values for the skyrmion radii and 10 values of the parameter , which after normalization changes the magnetization of the lattice. Since in our analysis we will be using snapshots from MC simulations as images, in this constructed dataset we do not require periodic boundary conditions, but we must take into account rotations and translations of the parametrization pattern, which we do considering 8 possible rotations and 3 different translations in the position along the 3 directions of the nearest neighbors bonds of the triangular lattice. Therefore, we have a training and validation dataset of different skyrmion crystals in triangular lattices. A few examples of the training snapshots with different skyrmion sizes are those illustrating the parametrisation in Fig. 1.
@@ III.2 Monte Carlo simulations dataset
We performed Monte Carlo Metropolis-Hastings simulations, combined with microcanonical updated (overrelaxation) on the three spin models described in the previous section: (1) TFDMxy: a ferromagnetic exchange and in plane DMI model in a triangular lattice under a magnetic field, fixing (2) KFDMxy: a similar model in a kagome lattice, for (3) KFDMz: the model presented in Eq. (3) including an out-of-plane Dzyaloshinskii-Moriya interaction . Simulations were done using periodic boundary conditions, and lowering the temperature from high temperature in 80 steps using the annealing technique. Up to 10 copies with independent seeds were done for each model and magnetic field. To apply the algorithms to the kagome snapshots, we chose one of the three sublattices (each one corresponding to the three types of sites in the kagome lattice, numbered in Table 1), so as to work with a snapshot.
@@ III.3 Algorithms
Here we give details on the two algorithms used in this work: Principal Component Analysis (PCA) and Convolutional Auto Encoder (CAE). In the Appendix we discuss the effect of choosing other parameters and architectures for these algorithms.
@@ III.3.2 Convolutional Autoencoder
In our work, we will use the “anomaly detection” approach. The idea is to train an AE for a given type of data, and then apply it to a larger dataset, comparing the input and the output with a convenient variable, such as the root mean square error. If the error is small, then the input and output are similar, if it is large, then the input data may not be of the same type as that used in training the AE, so a larger error signals an anomaly. As in the PCA, we will use the artificial skyrmion lattice dataset to train a convolutional autoencoder (CAE - an AE with a convolutional layer), measuring the RMSE. In this case, 80 of the data were used for training and 20 for validation. The RMSE for the trained CAE was 0.0178 in the training set and 0.0181 in the validation set.
The structure of the CAE is the following: the encoder consists of one convolutional layer with activation function ‘relu’, 32 filters, filter size 3 and padding ‘same’, followed by a MaxPooling layer with filter size 2 and padding ‘same’. The decoder consists of a convolutional layer, with the same characteristics as the first one, and an UpSampling layer. It finalizes with a convolutional layer with one filter of size 3, a sigmoid activation function and padding ‘same’. The CAE was trained and validated with batch size 64, using early stopping for a maximum of 300 epochs with a patience of 10. Implementation of the CAE was done in TensorFlow using Keras [68, 69].
@@ IV Results and discussion
In this section, we present the analysis of the comparison between the input and output data for the MC snapshots obtained when applying the PCA and CAE algorithms that were trained with the artificial set of skyrmion lattices. We start in Subsec. A focusing on the three ferromagnetic models introduced in Sec. II (Table 1). Then, we will explore whether the anomaly detection technique using the same skyrmion-lattice trained CAE is able to provide some insight in two different antiferromagnetic models, which we will describe in Subsec. B.
@@ IV.1 Application to ferromagnetic skyrmion models
First, we compare the RMSE obtained for both algorithms for the TFDMxy snapshots. In order for the method to work, we would expect the RMSE to be lower at the skyrmion lattice phase, which is found at intermediate magnetic fields and lower temperature, and higher in the other phases, such as the low field helical phase at lower magnetic fields, or the high temperature paramagnetic phase. In Fig. 3, we present a series of examples comparing the input Monte Carlo snapshot with the output (decoded) snapshot after applying trained PCA and CAE. The first five rows correspond to the different phases that arise at lower and intermediate temperatures as the magnetic field is increased: helices, bimerons, skyrmion lattice, skyrmion gas and ferromagnetic. At first glance, the decoded images seem very close to the input ones, both for PCA and CAE, even for the helical and the ferromagnetic cases, which are types of configurations that were not used to train the algorithms. In these examples, the RMSE (indicated in each decoded image) for PCA is even lower in the helical and ferromagnetic cases than in the skyrmion lattice. This is not ideal, since we are proposing to use a larger RMSE as an indicator that the system is further from an ordered skyrmion phase. However, there is a significant difference in the output images at higher temperatures, illustrated in the last row of Fig 3 for a snapshot in the paramagnetic phase: both the PCA and the CAE have a significantly higher RMSE.
@@
To further explore and quantify whether the RMSE obtained with PCA or CAE is a useful tool to distinguish between the low temperature phases, in Fig. 4, (top panel) we present the chirality from MC simulations, PCA and CAE RMSE and the BCE as a function of the external magnetic field at the lowest simulated temperature for the TFDMxy model (). The curves are averaged over 10 independent copies, and the error bars are calculated as the standard deviation. Notice that this is the same model that was presented in Fig. 2, and thus the chirality is the same as the one shown in the bottom panel of that figure. Let us recall that the chirality cannot be “learned” by the algorithms, since in this proposal only the spin component is used for training.
@@
Turning to the RMSE, comparing the RMSE calculated from the output images of both algorithms, it can be seen that the RMSE from PCA is in fact lowest for the ferromagnetic phase, highest for the skyrmion gas, and, within the error, does not significantly distinguish between the helical and skyrmion lattice phases. On the other hand, the RMSE obtained comparing the MC snapshots and the CAE decoded ones shows a more interesting behaviour, clearly distinguishing three regions: a low field one with higher RMSE, corresponding to the helical phase, an intermediate one with the lowest values (the skyrmion phases), and then a third region where the RMSE goes slightly but noticeably up with and flattens at the ferromagnetic phase. The BCE curve also seems to distinguish between these three regions, showing an abrupt change as the system enters the skyrmion phases, but it is lowest in the ferromagnetic phase, and not in the skyrmion region, where the CAE was trained. Therefore, given the behavior of the RMSE for different magnetic fields, we choose the RMSE from the CAE to apply the anomaly detection technique in other models. We emphasize that, regardless of the values of the obtained RMSEs and BCEs, in this work we are interested in using them to distinguish phases, and therefore we focus on the shape of the curves. As a comment, a closer inspection of the RMSE curve also suggests that the RMSE error may also be a useful variable, an idea we will shortly come back to.
To explore whether the overall behavior of these variables is independent of the ratio (and thus, from skyrmion size), in the range where skyrmions are stabilized, in the bottom panels of Fig. 4, we present the same analysis changing the values of DMI, in the TFDMxy Hamiltonian. A first observation is that the maximum value of the chirality density is smaller, simply because for smaller there are larger, and thus fewer, skyrmions for the same lattice size. We find that the behaviour of the variables obtained from CAE, RMSE and BCE, is qualitatively the same. For the RMSE from PCA, we see that it increases at lower fields for , since in this case, upon a small magnetic field the helices are easily distorted into bimeron type textures.
@@
Having chosen our variable, the RMSE between the MC input snapshot and the decoded output from the trained CAE, we now apply the CAE to the other two ferromagnetic skyrmion models, both in the kagome lattice with in-plane interactions (KFDMxy), and one with an additional out-of-plane DMI (KFDMz) (see Table 1). We plot the RMSE BCE and show the MC chirality density as a function of magnetic field at the lowest simulated temperature for both models in Fig. 5. Given that we are choosing one kagome sublattice to apply the CAE, which it is trained for triangular lattices, in the rest of this work the chirality presented for the kagome models is the sublattice chirality, calculated in the chosen triangular sublattice. Notice that the chirality obtained from MC simulations is very similar for both models in the kagome lattice, suggesting that the number of skyrmions and skyrmion size may be the same in both models, and gives no indication of exotic phases. However, this is not the case when exploring the RMSE curve. For the KFDMxy, left panel, the behavior is very similar to the triangular lattice case. However, for the KFDMz model, before flattening in the ferromagnetic phase, the RMSE goes clearly up, there is a sharp jump in the BCE, but, as already mentioned, no significant differences are seen in the chirality, which is the order parameter connected to the skyrmion phases. Thus, the RMSE suggests that there is a different phase at higher magnetic fields, for . Inspection of the snapshots reveals this is indeed the case; an example is shown in Fig. 6. It is an unusual high field bimeron + skyrmion phase, with possible glassy behaviour [40]. As mentioned in Sec.II, bimerons are usually found as intermediate textures between the helices and the formation of skyrmions, at lower magnetic fields. Regardless of the nature of the textures, the RMSE indicates that at higher fields the KFDMz presents a different phase than the other two ferromagnetic models. We compare the mean value and the standard deviation of the RMSE as a function of the external field (scaled to its saturation value in each model, ) for the three models in Fig. 7. Inspecting the mean value of the RMSE presented in the left panel, it is clear that the RMSE is quantitavely and qualitatevely very similar in the three models in three regions (corresponding to helices, skyrmions and ferromagnetic), but evidently there is an exotic phase for the KFDMz case before the system goes into the ferromagnetic phase. The jump of the RMSE for this model in this region is not seen in the other two cases, strongly supporting the idea that in that region of the system goes into a different type of ordering. Interestingly, this information may also be extracted from the standard deviation of the RMSE (), seen in the right panel of Fig. 7: in the TFDMxy and KFDMxy models, is significantly larger at small magnetic fields, goes down at intermediate fields and flattens to zero at higher fields. For KFDMz, goes up again before going to zero, signaling a different behavior in this model.